OpenStudy (anonymous):

HELP ME PLEASE!!!! How would I calculate the volume of the rectangular prism in cubic centimetres? The dimensions are 25mm length, 21mm width and height is 9mm. Originally the units were in millimetres.

OpenStudy (anonymous):

V = lwh

OpenStudy (anonymous):

V = (25)(21)(9)

OpenStudy (anonymous):

got it so times

OpenStudy (anonymous):

you have to multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so 4725

OpenStudy (anonymous):

V = 4725 mm ^3 but then it asks you in centimeters

OpenStudy (anonymous):

yea it does

OpenStudy (anonymous):

so you have to convert from mm^3 to cm^3

OpenStudy (anonymous):

so i write 4725cm3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then

OpenStudy (anonymous):

because it is 4725mm^3

OpenStudy (anonymous):

to convert 4725 from mm to cm. You have to divide it by 10

OpenStudy (anonymous):

Volume = length*width* height V = 25*21*9 then your result multiply by .001 to get cm^3

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

V=472.5 cm^3

OpenStudy (anonymous):

so okay

OpenStudy (anonymous):

it is because 10mm = 1cm so you divide it by 10

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

And how do I calculate the capacity

OpenStudy (anonymous):

got any idea?

OpenStudy (anonymous):

1sec

OpenStudy (anonymous):

How would I calculate the capacity of the rectangular prism in cubic centimetres? The dimensions are 25mm length, 21mm width and height is 9mm. In millimetres

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

You convert it to meters ^3 and multiply it by 1000 because 1m^3 = 1000Liters

OpenStudy (anonymous):

but what do i convert

OpenStudy (anonymous):

the answer in cm that you received to meters

OpenStudy (anonymous):

472.5cm3 in metres and then multiply by a 1000

OpenStudy (anonymous):

I made a mistake converting mm^3 to cm^3. Instead of diving it by 10. you divide it by 1000

OpenStudy (anonymous):

so V=4725mm^3 ---> 4725/1000 = 4.725cm^3

OpenStudy (anonymous):

so u talking about the volume

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

okay so now the capacity

OpenStudy (anonymous):

So now to get the capacity, first we need to convert the volume to meters.

OpenStudy (anonymous):

meters^3 *

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

4.725/1000000

OpenStudy (anonymous):

0.000004725

OpenStudy (anonymous):

is that metres

OpenStudy (anonymous):

meter ^3

OpenStudy (anonymous):

and then now to litres which is 0.0047250L right?

OpenStudy (anonymous):

0.004725L yes

OpenStudy (anonymous):

and thats the answer

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so altogether it is 4.725/1000000=0.000004725- this is metres

OpenStudy (anonymous):

what do i do to get litres

OpenStudy (anonymous):

1m^3 = 1000liter

OpenStudy (anonymous):

so multiply

OpenStudy (anonymous):

you have to convert it to meters^3 to get liters so multiply it by 1000

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Yes, first you have to find the volume. Then if you they ask for the capacity, convert it to m^3 and multiply it by 1000.

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

how do u know all this stuff?

OpenStudy (anonymous):

You can understand all the concepts very easily when you are in high lvl of math.

OpenStudy (anonymous):

I'm in year 8

OpenStudy (anonymous):

so u mean in like year 9,10,11 or 12

OpenStudy (anonymous):

I had Calculus BC in High School 12th(Senior)

OpenStudy (anonymous):

If you get to hit Calculus. Then Geometry/Algebra become very easy because you are constantly using it in Calculus in different ways.

OpenStudy (anonymous):

i still have alot of questions to ask u regarding my assigment, but i have to have dinner. You gonna be on?

OpenStudy (anonymous):

like 3 questions

OpenStudy (anonymous):

Sorry I can't It's 2:33am and I have to sleep for university tomorrow :(

OpenStudy (anonymous):

ohh u in america lol i'm in australia 6:35

OpenStudy (anonymous):

6:35pm here

OpenStudy (anonymous):

well i gtg have dinner bye

OpenStudy (anonymous):

see you, have fun.